# Automating Knowledge Discovery from Scientific Literature via LLMs: A Dual-Agent Approach with Progressive Ontology Prompting

### v1

@@ 1 Intervention Ontology and Literature
Source Number of Literature
Wiley Online Library 9,511
NIH PubMed 30,928
SpringerLink 6,549
GoogleScholar 8,662
ASHA 5,641
EBSCOHost 2,886
Table 3: Literature base for speech-language intervention discovery.
@@
1.2 Speech-language Therapy Literature Collection We cultivate a literature corpus within the domain of speech-language therapy as shown in Table 3 to facilitate the discovery of intervention methods. To conduct our literature search, we use a collection of carefully selected keywords drawn from a glossary of commonly used terms in speech-language therapy. These keywords include:“speech language therapy, speech language disorder, speech sound disorder, articulation disorder, speech intervention, language intervention, auditory discrimination, auditory processing disorder, phonological awareness, phonological processes, auditory perception, babbling, motor speech disorder, morpheme, phonology, prosody, stuttering, language impairment, speech language pathologist, speech and language therapist, babbling, expressive language delay, cleft speech disorder, autism spectrum disorder, developmental phonological disorder, developmental stuttering, phonological impairment, developmental dysarthria, down syndrome, swallowing disorder, communication impairment, articulation impairment, dyslexia, apraxia, dysarthria, dysphagia, communication disorder, expressive language disorder, dyspraxia, aphasia, augmentative and alternative communication, central auditory processing disorder, cleft lip and palate, down syndrome, fluency disorders, hearing loss, orofacial myofunctional disorders, spoken language disorders, written language disorders, acquired brain injury, apraxia of speech, auditory comprehension, literacy impairments, voice difficulties, language-based learning disabilities.”
2 Prompting Methods
2.1 POP Prompting
In the POP algorithm, the annotation prompt for a specific concept node is generated from a template $T$, which is constructed using the node’s context (the k-hop visited neighborhood) and the knowledge discoveries within that context. We carefully design the prompt to task LLM generate $T$ as follows:
Example:
Annotate: Participant
Prompts:
T1: Who is the participant in {#case study}?
Some example output of GPT3.5 for the above prompt is:
T1: What is the frequency used in {#case study}?
T3: {#Intervention}, which includes {#participant}, is studied in {#case study}. What frequency is used in this study?
Since template $T$ is composed of concept nodes visited prior to the current node, the final annotation prompt for the current node is derived by populating the knowledge discoveries associated with these concept nodes into $T$.
2.2 CoT and SelfRefine Prompting
@@
Instead of using an external LLM to provide evaluation feedback to explorer, we use CoT and SelfRefine prompting techniques as baselines to task explorer refine annotations independently. The prompts of CoT and SelfRefine are as follows:
@@
3 Annotation Example
Below is an example of an intervention annotation based on the study titled ’Intensive Treatment of Dysarthria Secondary to Stroke.’
@@ 3 Annotation Example
4 Speech-Language Intervention Knowledge Graph
Below is the speech-language intervention knowledge graph constructed using our knowledge discovery framework. We plan to make this knowledge base publicly accessible after our paper is published.
Figure 7: Graph database overview.

### v2

@@ Appendix B Speech-Language Intervention Annotation
B.1 Literature Corpus
We cultivate a literature corpus of 64,177 research articles within the domain of speech-language pathology to facilitate intervention discovery. To conduct our literature search, we use a collection of carefully selected keywords drawn from a glossary of commonly used terms in speech-language pathology. These keywords include:
@@ B.1 Literature Corpus
“speech language therapy, speech language disorder, speech sound disorder, articulation disorder, speech intervention, language intervention, auditory discrimination, auditory processing disorder, phonological awareness, phonological processes, auditory perception, babbling, motor speech disorder, morpheme, phonology, prosody, stuttering, language impairment, speech language pathologist, speech and language therapist, babbling, expressive language delay, cleft speech disorder, autism spectrum disorder, developmental phonological disorder, developmental stuttering, phonological impairment, developmental dysarthria, down syndrome, swallowing disorder, communication impairment, articulation impairment, dyslexia, apraxia, dysarthria, dysphagia, communication disorder, expressive language disorder, dyspraxia, aphasia, augmentative and alternative communication, central auditory processing disorder, cleft lip and palate, down syndrome, fluency disorders, hearing loss, orofacial myofunctional disorders, spoken language disorders, written language disorders, acquired brain injury, apraxia of speech, auditory comprehension, literacy impairments, voice difficulties, language-based learning disabilities.”
B.2 Annotation Ontology
Speech-language pathology provides interventions for individuals with speech-language deficits, improving their quality of life at various stages. When choosing an intervention, evidence-based practice (EBP) is attractive as it integrates research evidence from literature into the decision-making process to ensure high-quality patient care Law et al. [1996]. Research on interventions, especially those presenting clear frameworks and comprehensive case studies, serves as valuable guidance for designing EBPs. In this paper, we apply our automated framework to speech-language intervention discovery, with the intervention ontology shown in Figure 4. A detailed explanation of the concepts within the ontology is provided below:
  • Disorder represents the type of disorder that causes difficulties in an individual’s voice, speech, language, or swallowing functions.
  • Efficacy represents the conclusion about the effectiveness of the intervention.
  • Case Study represents a detailed examination of the intervention on a particular individual or group with communication disorders. The purpose of a case study is to provide a deep understanding of the patient’s unique needs and assess the intervention’s effectiveness.
  • Age represent the age of experiment participant or claimed target population of the intervention. The age is quantified with a granularity of half a year. We additionally convert age to age groups including “newborn, infants, toddlers, children, preschoolers, school-age children, older children, youth, teens, adolescents, adult, young adult, middle aged, aged, senior”. Each specific age may be associated with multiple age groups. For instance, an individual aged 13 years could be categorized into the ‘teens,’ ‘adolescents,’ and ‘children’ age groups.
  • Language represents the speaking language of the experiment participant in the case study of the intervention.
Appendix C CoT & Self-Refine Baselines
@@
Instead of using an external LLM to provide evaluation feedback to the explorer, we use CoT and SelfRefine prompting techniques as baselines to task explorer refine annotations independently. The prompts of CoT and SelfRefine are as follows:
@@
Appendix D Annotation Example
Below is an example of an intervention annotation from the paper ”Intensive Treatment of Dysarthria Secondary to Stroke” Mahler and Ramig [2012].
@@ Appendix D Annotation Example
Appendix E Speech-Language Intervention Knowledge Base
Through our intervention discovery framework, we constructed the first intervention knowledge graph in the speech-language pathology domain. Our intervention KG contains 33,148 nodes with 324,707 relations. We present some views of our intervention KG in Figure 7 and Figure 8.